OpenStudy (anonymous):

What is the common difference of the following sequence? 13, 27, 227, 247, …

OpenStudy (campbell_st):

27 - 13 = 14 227 - 27 = 200 247 - 227 = 20 there is no common difference

OpenStudy (anonymous):

thanks